If a 1.5 m long tube with a 0.5 cm opening has its first resonance point at 37 cm, what would be the theoretical location of L3 (in meters with 2 decimal places)? Is it possible to hear the third resonance point in the tube?
Answer #1

Given data l=1.5 m

1 m=100 cm

Here 1.5 m is given we have to convert to meters

So, 1.5 m*100=150 cm

Here from data length of first resonance L 1=37 cm

If it is a closed pipe then

    L 1: L 2:L 3=1:3:5:7:--------

L 3=5 L 1

=5(37)

L 3=185 cm

Here we got 185>150

:It is a closed pipe we can’t hear the third resonance

For open pipe L 1:L 2:L 3:=====1:2:3

L 3=3 L 1

    =3(37)

   =111 cm (111<150)

:Here we can listen the third resonance

i.e., L 3=111 cm